2. Suppose the company paid dividend of $26.99. That is, D0 is $26.99. What is the amount of dividend in Year 5? That is, what is D5? Assume that the dividends are expected to grow by 5% each year.

3. ABC,. Inc just paid a dividend of $13.81. The dividends are expected to grow by 23% in Years 1-3. After that, the dividends are expected to grow by 5% each year. If the required rate of return is 17%, what is today's price of the stock?
Answer #1

Dividend in year 5 = D0(1+growth rate)^5

= 26.99(1.05)^5

= $34.45

3.Price of stock is equal to present value of all future dividends

=13.81(1.23)/(1.17) + 13.81(1.23)2/(1.17)2 + 13.81(1.23)3/(1.17)3 + 13.81(1.23)3(1.05)/(1.17)3(17%-5%)

= $186.22